On January 2, 2019, Collins Company purchased a delivery truck for $45,000 cash. The truck had an estimated useful life of seven years and an estimated residual value of $3,000. Straight-line depreciation was used. Assuming the transactions have commercial substance, prepare the journal entries to record the disposition of the truck on September 1, 2019, under each of the following assumptions: (20 marks)
(a) The truck and $55,000 cash were exchanged for equipment that had a fair value of $70,000.
(b) The truck and $40,000 cash were exchanged for a new delivery truck that had a fair value of
$70,000.
(c) If you were a new accountant in a company and you were asked to determine how best to depreciate a new machine they just bought, how would you determine the best method? What information would be useful to you in determining the best method?
Answer #1

А B 1 Depreciation for year=[cost-salvage]/life =[45000-30001/7 = 42000/7 = $ 6000 4 - Debit (S) Credit (S) 70000 4000 26000